# Detection of pairwise kSZ effect with DESI galaxy groups and Planck in Fourier space

### 07 Jan 2024 v1

@@ 4.6 Covariance matrix estimation
We perform the delete-one jackknife (JK) method (Sugiyama et al., 2018a) to estimate the data covariance through resampling of the date itself. We employ the kmeans algorithm [^9] https://github.com/esheldon/kmeans_radec/ on the random catalog to divide the group samples into $N_{\rm JK}$ spatial sub-regions on the sky (Kwan et al., 2017). We employ $N_{\rm JK}=200$, in which $N_{\rm JK}^{(\rm NGC)}=105$ and $N_{\rm JK}^{(\rm SGC)}=95$ so that their ratio is close to corresponding sky area ratio.
@@
to get an unbiased precision matrix $\hat{C}^{-1}$. Although the original derivation of the Hartlap factor is not based on the Jackknife technique, we apply this factor here to calculate a conservative precision matrix for our parameter inference procedure.
@@ 4.7 Survey window functions
Figure 4: Left: Multipoles of the survey window function for the baseline sample, which are used to compute the masked theoretical pairwise kSZ power dipole. Solid and dashed lines indicate the monopole and qradrupole separately. The red and blue lines represent the window functions of NGC and SGC respectively. Right: Masked (solid lines) and unmasked (dashed lines) theoretical predictions of the pairwise kSZ dipole for the baseline sample. The red and blue lines are for NGC and SGC respectively. On the lower panel, the relative difference ${\rm Diff}=P_{\rm mask}/P_{\rm nomask}-1$ induced by the window function effect is presented. The inflation of ‘Diff’ at $k\sim 0.15h/\rm Mpc$ is due to the zero-crossing of $P^{\ell=1}_{\rm kSZ}$ at that scale.
@@
where $RR$ is the number of random pairs in a logarithmic $s$ bin [^10] Corrfunc.mocks.DDsmu_mocks. Here $Q_{\ell}$ is normalized to enforce $Q_{0}(0)=1$. In order to speed up the calculation, for samples with $M_{\rm th}\geq 10^{13}{\rm M}_{\odot}/h$, we randomly dilute the random catalog to the size of the $M_{\rm th}=10^{13}{\rm M}_{\odot}/h$ group sample, and for samples with $M_{\rm th}<10^{13}{\rm M}_{\odot}/h$, we dilute the random catalog to the size of the corresponding group sample size. A convergence test in appendix B shows that this dilution induces negligible systematic effects on the calculation of $\tilde{P}_{\rm kSZ}^{\ell=1}$ in eq. (43).
The left panel of figure 4 shows the multipole components of the survey window function applied in our analysis. We expect that monopole and quadrupole respectively tend to unit and zero when $s\rightarrow{}0$, and both of them are equal to zero in dimensions beyond the survey volume. The positive quadrupole values reflect that for a fixed $s$ bin, there are more randoms pairs found along the LOS than perpendicular to the LOS, and vice versa.
In turn, the masked kSZ power spectrum including the window function can be theoretically calculated by Sugiyama et al. (2018a)
@@
Here, $j_{\ell}$ is the spherical Bessel function of order $\ell$. For the dipole, the expansion of eq. (41) can be linearly truncated as
@@
in which the contribution of higher multipoles of the kSZ correlation function (e.g. the octopole $\xi_{\rm kSZ}^{\ell=3}$) are small and can be ignored. The robustness of this linear truncation is verified in appendix C. Eq. (43) is the final expression used to fit the measurement of the pairwise kSZ power dipole in this work. Its detailed derivation is presented in appendix C of Sugiyama et al. (2018a).
The theoretical predictions of the pairwise kSZ dipoles with and without the survey window functions are shown on the right panel of figure 4. The amplitude of the masked $P^{\ell=1}_{\rm kSZ}$ becomes smaller than the true one at large scales, due to the lack of group pairs beyond the survey volume. The window function effect becomes significant at $k<0.05h/\rm Mpc$ and is of the order of $10-20\%$ around $k\sim 0.01h/\rm Mpc$. Since our measurement S/N is dominated by linear scales, this effect is a key systematic error that ought to be corrected in our analysis.
@@ 4.8 Imaging systematics
Figure 5: Left: The pairwise kSZ dipole measurement of the baseline sample without (data points with error bars) and with (dashed line) the imaging-systematics correction. Right: The multipoles of redshift space density power spectrum without (data points with error bars) and with (dashed line) the imaging-systematics correction. The lower panel shows the fractional difference between those of “corrected” and “raw” ones. Measurements are made from teh baseline sample.
@@
It has been well demonstrated that a variety of observation conditions, such as stellar contamination, Galactic extinction, sky brightness, seeing, and airmass, can introduce spurious fluctuations in the measured galaxy density field (e.g. Scranton et al., 2002; Myers et al., 2006; Morrison & Hildebrandt, 2015; Kitanidis et al., 2020). These fluctuations, known as imaging systematics, must be addressed in order to obtain reliable results for large-scale clustering studies. The group catalogs used in this work are derived from photometric galaxy catalogs, which may be subject to such biases.
@@
A widely-used approach to reduce the imaging systematics is to assign weights to the targets so that the weighted sample does not show any correlation with the various observation conditions (e.g. Ross et al., 2017; Bautista et al., 2018). This can be achieved by minimizing the difference between the predicted and observed target surface density. The predicted target density is simply a function of various imaging maps, which are designed to not contain any cosmological signal. Previous studies assumed a linear functional form (e.g. Ross et al., 2011), which may not be able to capture the non-linear dependence on imaging systematics in areas with strong contamination, such as close to the Galactic plane (e.g. Ho et al., 2012). Recently, a machine learning-based approach has been used to capture these complex relationships (e.g. Rezaie et al., 2020; Chaussidon et al., 2022).
@@
We reduce the imaging systematics in our targets using the Random Forest algorithm, which is available in the GitHub code regressis [^11] https://github.com/echaussidon/regressis developed by Chaussidon et al. (2022). Xu et al. (2023) adapted the code to remove imaging systematics in auto- and cross-correlations between multiple tomography bins in the galaxy samples from Data Release 9 of the DESI Legacy Imaging Survey. Here, we follow the steps outlined in Xu et al. (2023) to mitigate imaging systematics and refer interested readers to their paper for details.
@@
The comparison between the measured $\hat{P}_{\rm kSZ}^{\ell}$ with and without correcting imaging systematics is shown on the left panel of figure 5. We see little systematic offsets when we ignore the imaging systematics. On the right panel of figure 5, we also find little differences between redshift space density power spectrum multipoles $\hat{P}_{\ell}$ with and without the imaging systematics correction, in which $\hat{P}_{\ell}$ is analyzed in appendix A to constrain free parameters and determine the theoretical $P_{\rm pv}$ template in the $\bar{\tau}$ fitting. As a result, we consider the imaging systematics as a negligible systematic error source in our analysis and we will not correct them in this work.
@@ 5 Results
In this section we presents the main results of this paper, including the pairwise kSZ dipoles, the fitted $\bar{\tau}$ values and the $\bar{\tau}-M$ relationship.
@@ 5.1 Fitting procedure
We estimate the average optical depth $\bar{\tau}$ of each group sample by fitting the measured kSZ dipole with the template given by eq. (43). $\bar{\tau}$ is the only free parameter to be fitted. The $\chi^{2}$ of the fitting is
@@ 5.2 Pairwise kSZ dipole and $\bar{\tau}$
Figure 7: Left: Measured $\bar{\tau}$ profiles of the baseline sample by choosing different AP filter radii ($\theta_{\rm AP}$). The blue dots with error bars represent the $\bar{\tau}$ profile of the baseline sample, and the black dots with error bars show the results of the null test by targeting a random sample having the same size of the baseline sample. On top of the figure we label the comoving radius corresponding to $\theta_{\rm AP}$. Right: Corresponding detection $\rm S/N$s.
@@
The left panel of figure 6 is the main result of this work, which shows the highest ${\rm S/N}=5.2$ dipole measurement from the baseline sample with $\theta_{\rm AP}=4.2^{\prime}$ and its covariance. This serves as one of the highest S/N kSZ detections in the literature, and is the highest S/N kSZ detection in Fourier space currently. As expected, the useful kSZ signal dominates only at linear scales since it is diluted by the photo-z damping and noises (shot noise, CMB detector noise and primary CMB contamination) at non-linear scales. As a result the linear kSZ model (solid lines) gives a good fit to the data, and the fitted $\bar{\tau}=(3.96\pm 0.76)\times 10^{-5}$ for $\tilde{M}=1.75\times 10^{13}{\rm M}_{\odot}/h$. On the right panel of figure 6 we present the correlation coefficients of the dipole measurements from baseline sample. Off-diagonal elements are sub-dominant in the covariance matrix and the kSZ field on each $k$ scale evolves independently with each other.
By varying $\theta_{\rm AP}$, we display in figure 7 the measured $\bar{\tau}$ profile of the baseline sample. This profile encodes valuable information about the gas distribution within halos, and it can be fitted by some gas density profile model to constrain the gas fraction $f_{\rm gas}$ within halos (Amodeo et al., 2021; Sugiyama et al., 2018a). In a companion paper we will study this $\bar{\tau}$ profile with adequate profile models, e.g. the baryonification models (Schneider & Teyssier, 2015; Giri & Schneider, 2021; Aricò et al., 2021), and constrain the halo gas content in our samples.
@@ 5.3 $\bar{\tau}-\tilde{M}$ relationship
Figure 8: The measured $\bar{\tau}-\tilde{M}$ relationship with $\theta_{\rm AP}=4.2^{\prime}$. The points with error bars are measurements from various samples with different $M_{\rm th}$ and the solid lines are the best fitted eq. (48) with $\beta=(0.98^{+0.15}_{-0.15})\times 10^{-4}$ and $\gamma=0.55^{+0.10}_{-0.10}$. We caution that cross-correlations between different $\bar{\tau}$s are ignored in this fitting, so the fitted error bars for $\beta$ and $\gamma$ are underestimated here.
Next we explore the relationship between the average optical depth and the median halo mass of a sample. We choose the measured optical depths of different samples in appendix D measured with $\theta_{\rm AP}=4.2^{\prime}$ and fit the $\bar{\tau}-\tilde{M}$ relationship with a simple power-law,
@@
The results are shown in figure 8. Red data points with error bars are measured $\bar{\tau}$s associated with the median group sample masses and the red solid line is the best fitted eq. (48). The fitted $\gamma=0.55^{+0.1}_{-0.1}$ is roughly consistent with the $\bar{\tau}\propto\tilde{M}$ relation, as expected for massive groups, and this result confirms our motivation of sample selection with the group mass. In Chen et al. (2022), a higher value of $\gamma=1.25\pm 0.14$ is found. Although being deviated from each other, the discrepancy between two results is not as large as it looks here. Since group samples in both works are not independent of each other by definition, the uncertainty of $\gamma$ is moderately underestimated in both papers. Albeit this, the disagreement can come from several factors. (1) A minor one is that two works use different AP filter radii, with $\theta_{\rm AP}=3.4^{\prime}$ in Chen et al. (2022) and $\theta_{\rm AP}=4.2^{\prime}$ here. (2) In Chen et al. (2022), most group samples used to fit eq. (48) are heavy clusters with masses larger than $10^{14}{\rm M}_{\odot}/h$, while in our work the mass range of group samples is ranging from $10^{13}{\rm M}_{\odot}/h$ to $10^{14}{\rm M}_{\odot}/h$. (3) The redshift distribution of the group sample is affected by the mass threshold, so the sample redshift distributions are not identical to each other, and can introduce potential systematic biases in the measured $\bar{\tau}-\tilde{M}$ relationship. (4) Two works treat the density biases ($b_{1}$, $b_{2}$) and the photo-z uncertainty $\sigma_{\rm pho-z}$ in different manners. In this work, we apply a joint analysis of kSZ and RSD effects on the same dataset. We first constrain $b_{1}$, $b_{2}$ and $\sigma_{\rm pho-z}$ from the measured $\hat{P}_{\ell}$, these determined free parameters are then adopted to determine $P_{\rm pv}$ and fit $\bar{\tau}$. In Chen et al. (2022), they evaluated $b_{1}$ from the Sheth01 $b_{1}-M$ model (Sheth et al., 2001). Then $\xi_{\rm pv}$ was directly measured from the simulation halo catalog constructed to match the calculated $b_{1}$, on which the photo-z uncertainties are directly applied on each halo.
@@
In Chen et al. (2022), a fiducial sample of the richest $1.2\times 10^{5}$ groups with $\log\bar{M}=14.19$ was constructed from a previous version of our adopted group catalog, which is based on the Data Released 8 (DR8) of the the DESI Legacy Imaging Survey. The pairwise kSZ correlation function was measured from the fiducial sample and the optical depth is fitted to be $\bar{\tau}=1.66\pm 0.35\times 10^{-4}$, namely a $\sim 5\sigma$ detection as well. We confirm that this measurement is consistent with our fitting result regarding eq. (48) within errors, and this is a good cross-check of the kSZ data analysis in both configuration and Fourier spaces [^12] Here we ignore difference between mean and median masses, and also ignore the $\theta_{\rm AP}$ difference between $\theta_{\rm AP}=3.4^{\prime}$ in Chen et al. (2022) and our $\theta_{\rm AP}=4.2^{\prime}$, since the $\bar{\tau}$ variation is not significant from $\theta_{\rm AP}=3.4^{\prime}$ to $4.2^{\prime}$ for heavy groups, as shown in figure 14..

### 01 Mar 2024 v2

@@ 4.6 Covariance matrix estimation
We perform the delete-one jackknife (JK) method (Sugiyama et al., 2018a) to estimate the data covariance through resampling of the date itself. We employ the kmeans algorithm [^9] https://github.com/esheldon/kmeans_radec/ on the random catalog to divide the group samples into $N_{\rm JK}$ spatial subregions on the sky (Kwan et al., 2017). We employ $N_{\rm JK}=200$, in which $N_{\rm JK}^{(\rm NGC)}=105$, and $N_{\rm JK}^{(\rm SGC)}=95$ so that their ratio is close to corresponding sky area ratio.
@@
to get an unbiased precision matrix $\hat{C}^{-1}$. Although the original derivation of the Hartlap factor is not based on the JK technique, we apply this factor here to calculate a conservative precision matrix for our parameter inference procedure.
@@ 4.7 Survey window functions
Figure 4: Left: Multipoles of the survey window function for the baseline sample, which are used to compute the masked theoretical pairwise kSZ power dipole. Solid and dashed lines indicate the monopole and qradrupole separately. The red and blue lines represent the window functions of NGC and SGC respectively. Right: Masked (solid lines) and unmasked (dashed lines) theoretical predictions of the pairwise kSZ dipole for the baseline sample. The red and blue lines are for NGC and SGC respectively. On the lower panel, the relative difference ${\rm Diff}=P_{\rm mask}/P_{\rm nomask}-1$ induced by the window function effect is presented. The inflation of “Diff” at $k\sim 0.15h/\rm Mpc$ is due to the zero-crossing of $P^{\ell=1}_{\rm kSZ}$ at that scale.
@@
where $RR$ is the number of random pairs in a logarithmic $s$ bin [^10] Corrfunc.mocks.DDsmu_mocks. Here, $Q_{\ell}$ is normalized to enforce $Q_{0}(0)=1$. In order to speed up the calculation, for samples with $M_{\rm th}\geq 10^{13}M_{\odot}/h$, we randomly dilute the random catalog to the size of the $M_{\rm th}=10^{13}M_{\odot}/h$ group sample, and for samples with $M_{\rm th}<10^{13}M_{\odot}/h$, we dilute the random catalog to the size of the corresponding group sample size. A convergence test in Appendix B shows that this dilution induces negligible systematic effects on the calculation of $\tilde{P}_{\rm kSZ}^{\ell=1}$ in Equation (43).
The left panel of figure 4 shows the multipole components of the survey window function applied in our analysis. We expect that monopole and quadrupole respectively tend to unit and zero when $s\rightarrow{}0$, and both of them are equal to zero in dimensions beyond the survey volume. The positive quadrupole values reflect that, for a fixed $s$ bin, there are more random pairs found along the LOS than perpendicular to the LOS, and vice versa.
In turn, the masked kSZ power spectrum including the window function can be theoretically calculated by (Sugiyama et al., 2018a)
@@
Here, $j_{\ell}$ is the spherical Bessel function of order $\ell$. For the dipole, the expansion of Equation (41) can be linearly truncated as
@@
in which the contributions of higher multipoles of the kSZ correlation function (e.g., the octopole $\xi_{\rm kSZ}^{\ell=3}$) are small and can be ignored. The robustness of this linear truncation is verified in Appendix C. Equation (43) is the final expression used to fit the measurement of the pairwise kSZ power dipole in this work. Its detailed derivation is presented in Appendix C of Sugiyama et al. (2018a).
The theoretical predictions of the pairwise kSZ dipoles with and without the survey window functions are shown on the right panel of figure 4. The amplitude of the masked $P^{\ell=1}_{\rm kSZ}$ becomes smaller than the true one at large-scales, due to the lack of group pairs beyond the survey volume. The window function effect becomes significant at $k<0.05h/\rm Mpc$ and is of the order of $10\%-20\%$ around $k\sim 0.01h/\rm Mpc$. Since our measurement S/N is dominated by linear scales, this effect is a key systematic error that ought to be corrected in our analysis.
@@ 4.8 Imaging systematics
Figure 5: Left: the pairwise kSZ dipole measurement of the baseline sample without (data points with error bars) and with (dashed line) the imaging-systematics correction. Right: The multipoles of redshift space density power spectrum without (data points with error bars) and with (dashed line) the imaging-systematics correction. The lower panel shows the fractional difference between those of “corrected” and “raw” ones. Measurements are made from the baseline sample.
@@
It has been well demonstrated that a variety of observation conditions, such as stellar contamination, Galactic extinction, sky brightness, seeing, and airmass, can introduce spurious fluctuations in the measured galaxy density field (e.g., Scranton et al., 2002; Myers et al., 2006; Morrison & Hildebrandt, 2015; Kitanidis et al., 2020). These fluctuations, known as imaging systematics, must be addressed in order to obtain reliable results for large-scale clustering studies. The group catalogs used in this work are derived from photometric galaxy catalogs, which may be subject to such biases.
@@
A widely used approach to reduce the imaging systematics is to assign weights to the targets so that the weighted sample does not show any correlation with the various observation conditions (e.g. Ross et al., 2017; Bautista et al., 2018). This can be achieved by minimizing the difference between the predicted and observed target surface density. The predicted target density is simply a function of various imaging maps, which are designed to not contain any cosmological signal. Previous studies assumed a linear functional form (e.g., Ross et al., 2011), which may not be able to capture the nonlinear dependence on imaging systematics in areas with strong contamination, such as close to the Galactic plane (e.g., Ho et al., 2012). Recently, a machine-learning-based approach has been used to capture these complex relationships (e.g., Rezaie et al., 2020; Chaussidon et al., 2022).
@@
We reduce the imaging systematics in our targets using the random forest algorithm, which is available in the GitHub code regressis [^11] https://github.com/echaussidon/regressis developed by Chaussidon et al. (2022). Xu et al. (2023) adapted the code to remove imaging systematics in autocorrelations and cross-correlations between multiple tomography bins in the galaxy samples from DR9 of the DESI Legacy Imaging Survey. Here, we follow the steps outlined in Xu et al. (2023) to mitigate imaging systematics and refer interested readers to their paper for details.
@@
The comparison between the measured $\hat{P}_{\rm kSZ}^{\ell}$ with and without correcting imaging systematics is shown on the left panel of figure 5. We see little systematic offsets when we ignore the imaging systematics. On the right panel of figure 5, we also find little differences between redshift space density power spectrum multipoles $\hat{P}_{\ell}$ with and without the imaging-systematics correction, in which $\hat{P}_{\ell}$ is analyzed in Appendix A to constrain free parameters and determine the theoretical $P_{\rm pv}$ template in the $\bar{\tau}$ fitting. As a result, we consider the imaging systematics as a negligible systematic error source in our analysis, and we will not correct them in this work.
@@ 5 Results
In this section, we presents main results of this paper, including the pairwise kSZ dipoles, the fitted $\bar{\tau}$ values, and the $\bar{\tau}-\tilde{M}$ relationship.
@@ 5.1 Fitting procedure
We estimate the average optical depth $\bar{\tau}$ of each group sample by fitting the measured kSZ dipole with the template given by Equation (43). $\bar{\tau}$ is the only free parameter to be fitted. The $\chi^{2}$ of the fitting is
@@ 5.2 Pairwise kSZ dipole and $\bar{\tau}$
Figure 7: Left: Measured $\bar{\tau}$ profiles of the baseline sample by choosing different AP filter radii ($\theta_{\rm AP}$). The blue dots with error bars represent the $\bar{\tau}$ profile of the baseline sample, and the black dots with error bars show the results of the null test by targeting a random sample having the same size of the baseline sample. On top of the figure, we label the comoving radius corresponding to $\theta_{\rm AP}$. Right: corresponding detection $\rm S/N$s.
@@
The left panel of figure 6 is the main result of this work, which shows the highest ${\rm S/N}=5.2$ dipole measurement from the baseline sample, with $\theta_{\rm AP}=4.2^{\prime}$ and its covariance. This serves as one of the highest S/N kSZ detections in the literature, and is the highest S/N kSZ detection in Fourier space currently. As expected, the useful kSZ signal dominates only at linear scales since it is diluted by the photo-z damping and noises (shot noise, CMB detector noise, and primary CMB contamination) at nonlinear scales. The nonlinear kSZ model (solid lines) gives a good fit to the data, and the fitted $\bar{\tau}=(3.96\pm 0.76)\times 10^{-5}$ for $\tilde{M}=1.75\times 10^{13}M_{\odot}/h$. On the right panel of figure 6 we present the correlation coefficients of the dipole measurements from baseline sample. The off-diagonal elements are subdominant in the covariance matrix, and the kSZ field on each $k$ scale evolves independently with each other.
By varying $\theta_{\rm AP}$, we display in figure 7 the measured $\bar{\tau}$ profile of the baseline sample. This profile encodes valuable information about the gas distribution within halos, and it can be fitted by some gas density profile model to constrain the gas fraction $f_{\rm gas}$ within halos (Amodeo et al., 2021; Sugiyama et al., 2018a). In a companion paper, we will study this $\bar{\tau}$ profile with adequate profile models, e.g., the baryonification models (Schneider & Teyssier, 2015; Giri & Schneider, 2021; Aricò et al., 2021), and constrain the halo gas content in our samples.
@@ 5.3 $\bar{\tau}-\tilde{M}$ relationship
Figure 8: The measured $\bar{\tau}-\tilde{M}$ relationship with $\theta_{\rm AP}=4.2^{\prime}$. The points with error bars are measurements from various samples with different $M_{\rm th}$ and the solid lines are the best-fitted Equation (48) with $\beta=(0.98^{+0.15}_{-0.15})\times 10^{-4}$ and $\gamma=0.55^{+0.10}_{-0.10}$. We caution that cross-correlations between different $\bar{\tau}$s are ignored in this fitting, so the fitted error bars for $\beta$ and $\gamma$ are underestimated here.
Next, we explore the relationship between the average optical depth and the median halo mass of a sample. We choose the measured optical depths of different samples in appendix D measured with $\theta_{\rm AP}=4.2^{\prime}$ and fit the $\bar{\tau}-\tilde{M}$ relationship with a simple power law,
@@
The results are shown in figure 8. The red data points with error bars are measured $\bar{\tau}$s associated with the median group sample masses, and the red solid line is the best-fitted Equation (48). The fitted $\gamma=0.55^{+0.1}_{-0.1}$ is roughly consistent with the $\bar{\tau}\propto\tilde{M}$ relation, as expected for massive groups, and this result confirms our motivation of sample selection with the group mass. In Chen et al. (2022), a higher value of $\gamma=1.25\pm 0.14$ is found. Although being deviated from each other, the discrepancy between two results is not as large as it looks here. Since group samples in both works are not independent of each other by definition, the uncertainty of $\gamma$ is moderately underestimated in both papers. Albeit, with this, the disagreement can come from several factors. (1) A minor one is that two works use different AP filter radii, with $\theta_{\rm AP}=3.4^{\prime}$ in Chen et al. (2022), and $\theta_{\rm AP}=4.2^{\prime}$ here. (2) In Chen et al. (2022), most group samples used to fit Equation (48) are heavy clusters with masses larger than $10^{14}{\rm M}_{\odot}/h$, while in our work the mass range of group samples is ranging from $10^{13}{\rm M}_{\odot}/h$ to $10^{14}{\rm M}_{\odot}/h$. (3) The redshift distribution of the group sample is affected by the mass threshold, so the sample redshift distributions are not identical to each other, and can introduce potential systematic biases in the measured $\bar{\tau}-\tilde{M}$ relationship. (4) Two works treat the density biases ($b_{1}$, $b_{2}$) and the photo-z uncertainty $\sigma_{\rm pho-z}$ in different manners. In this work, we apply a joint analysis of kSZ and RSD effects on the same data set. We first constrain $b_{1}$, $b_{2}$ and $\sigma_{\rm pho-z}$ from the measured $\hat{P}_{\ell}$, these determined free parameters are then adopted to determine $P_{\rm pv}$ and fit $\bar{\tau}$. In Chen et al. (2022), they evaluated $b_{1}$ from the Sheth01 $b_{1}-M$ model (Sheth et al., 2001). Then, $\xi_{\rm pv}$ was directly measured from the simulation halo catalog constructed to match the calculated $b_{1}$, on which the photo-z uncertainties are directly applied on each halo.
@@
In Chen et al. (2022), a fiducial sample of the richest $1.2\times 10^{5}$ groups with $\log\bar{M}=14.19$ was constructed from a previous version of our adopted group catalog, which is based on the Data Release 8 of the the DESI Legacy Imaging Survey. The pairwise kSZ correlation function was measured from the fiducial sample, and the optical depth is fitted to be $\bar{\tau}=1.66\pm 0.35\times 10^{-4}$, namely a $\sim 5\sigma$ detection as well. We confirm that this measurement is consistent with our fitting result regarding Equation (48) within errors, and this is a good cross-check of the kSZ data analysis in both configuration and Fourier spaces [^12] Here, we ignore difference between mean and median masses, and also ignore the $\theta_{\rm AP}$ difference between $\theta_{\rm AP}=3.4^{\prime}$ in Chen et al. (2022) and our $\theta_{\rm AP}=4.2^{\prime}$, since the $\bar{\tau}$ variation is not significant from $\theta_{\rm AP}=3.4^{\prime}$ to $4.2^{\prime}$ for heavy groups, as shown in figure 14..